# Hierarchical Context Alignment with Disentangled Geometric and Temporal Modeling for Semantic Occupancy Prediction

### v1

@@ 3.1.1 Preliminary
where $\widehat{\mathbf{V}}$ represents the estimated 3D semantic voxel grid, which aims to approximate the ground truth semantic occupancy or LiDAR semantic labels. For LiDAR semantic segmentation, we use the LiDAR data only for point query to compute evaluation metrics following previous works [6, 17].
3.1.2 Architectural Design Comparison and Analysis
To estimate high-quality 3D semantic voxel grid $\widehat{\mathbf{V}}$, existing methods attempt to optimize the scene modeling process from geometric or temporal perspectives while neglecting the misalignment issue. Specifically, the geometric modeling solution [17] leverages naive monocular depth estimation for semantic voxel grid estimation and neglects the uncertainty inherent in the depth estimation process:
@@ 3.1.2 Architectural Design Comparison and Analysis
where $\otimes$ denotes the outer product. Extr and Mono represent the 2D feature extractor and monocular depth estimator, respectively. Such a process inevitably causes geometric ambiguity when the estimated depth is integrated with the corresponding 2D features. On the other hand, the temporal modeling solution [5] straightforwardly stacks the temporal image frames to construct the semantic voxel grid $\widehat{\mathbf{V}}$, which ignores the positional changes of shared semantic content across different perspectives:
@@
To tackle the misalignment issue, we explore first disentangling the complicated semantic scene comprehension into geometric and temporal context alignment, and further align these contexts globally to compose them together to construct reliable semantic voxel grid $\widehat{\mathbf{V}}$:
$\widehat{\mathbf{V}}=\texttt{Compose}(\texttt{Geo}(I_{t}^{{rgb}},I_{t-1}^{{rgb}},\ldots),\texttt{Tem}(I_{t}^{{rgb}},I_{t-1}^{{rgb}},\ldots)).$ (4)
In this way, our methods take advantage of the complementary merits of the geometric and temporal representations in a disentangled local-global architecture, which are hierarchically aligned for a reliable context composition.
@@ 3.1.3 Overall Framework
Geometric Alignment. The voxel feature volume $\mathbf{V}_{{vox}}$ is constructed using a UNet architecture based on a pre-trained EfficientNetB7 [64]. The network initially generates features with spatial dimensions of $\mathbb{R}^{H/4\times W/4}$. We then adopt the lifting process following previous studies [30, 37, 18], to form $\mathbf{V}_{{vox}}$ from the contextual information and depth distribution. For depth distribution modeling, we use off-the-shelf monocular [65] or stereo [66] depth estimation networks. By default, we leverage stereo depth estimation to form our stereo-based pipeline of Hi-SOP(S). Additionally, a monocular-based pipeline of Hi-SOP(M) is presented to enhance versatility for scenarios lacking stereo inputs. To facilitate reliable geometry alignment when constructing $\mathbf{V}_{{vox}}$, we employ a Geometric Confidence-awareness Lifting (GCL) module, which is detailed in Section 3.2.
@@
Temporal Alignment. To construct the temporal feature volume $\mathbf{V}_{{tem}}$, we feed current and historical frames into a lightweight PoseNet [67, 59] to generate temporal feature volume $\mathbf{V}_{{tem}}$ using homography warping. Different from computing matching costs in typical temporal depth estimation methods [58, 57, 59], we aim to preserve context features within $\mathbf{V}_{{tem}}$. The details are presented in Section 3.3.
@@
Following that, we leverage the temporal volume $\mathbf{V}_{{tem}}$ to generate the cross-frame affinity $\hat{\mathbf{A}}$, quantifying contextual relevance/correspondences between current and historical features. This affinity is then used to reassemble the temporal content and dynamically refine the sampling locations, resulting in a reliable aligned temporal volume $\widetilde{\mathbf{V}}_{{tem}}$. Further details on Cross-frame Pattern Affinity (CPA) and Affinity-based Dynamic Refinement (ADR) are presented in Sections 3.4 and 3.5, respectively.
@@
Global Composition. To construct a reliable unified representation with semantically consistentency, the temporal feature volume $\widetilde{\mathbf{V}}_{{tem}}$ is aligned with the voxel feature volume $\mathbf{V}_{{vox}}$ in a global view through depth-hypothesis based transformation. These two volumes are composed together to generate the composed volume $\mathbf{V}_{{com}}$. The Depth-Hypothesis-Based Transformation (DHBT) module is detailed in Section 3.6.
3.2 Geometric Alignment with Confidence-aware Lifting
@@
As introduced in previous works [30, 18], the lifting process establishes volumetric features that store pixel-level context with their associated depth distribution. Nonetheless, [18] highlights that this process is inherently ambiguous and prone to producing unreliable representations in challenging regions (e.g., severe occlusion and high reflection) where the depth estimation results are unreliable [18]. To facilitate reliable geometry modeling within the voxel feature volume $\mathbf{V}_{{vox}}$, we propose a Geometric Confidence-aware Lifting (GCL) module to explicitly model the geometric information with depth distribution confidence.
@@
As shown in Figure 4, the module takes the depth feature $\textbf{F}_{d}$ from the depth net and the context feature $\textbf{F}_{c}$ from the context net as inputs. To establish pixel-level reliable information for dense prediction, we develop a depth confidence-aware cross-attention mechanism to explicitly indicate the confidence information of the depth distribution and take advantage of the relevant context to complement the low-confidence regions. Specifically, to project $\textbf{F}_{d}$ to a confidence map $\textbf{C}_{d}$ , we first adopt $softmax$ to convert depth cost value $d_{i}$ of $\textbf{F}_{d}$ into a probability form, and then take out the highest probability value among all depth hypothesis planes along the depth dimension as the prediction confidence. The process is formally written as:
@@
where the $softmax$ is applied across the depth dimension and $\mathrm{WTA}$ represents winner-takes-all operation. $D_{max}$ denotes the length of the depth dimension.
Next, we utilize the depth confidence information to enforce the cross-attention for pixel-level reliable geometric modeling. Specifically, we obtain the query $Q_{d}$ from $\textbf{F}_{d}$ by flattening in spatial and depth dimensions following standard protocol [68, 69]. Similarly, the context feature $\textbf{F}_{c}$ is forwarded and its key and value are denoted as $K_{c},V_{c}$, respectively. To reduce computational and memory consumption, we follow [70, 71] to compute linear cross-attention:
@@
3.3 Temporal Alignment with Feature Volume Construction
The fine-grained nature of the SOP task requires constructing temporally aligned features for accurate and robust perception. Instead of simply stacking input images from various viewpoints as [5], we propose to align the temporal invariant content using explicit homography transformation.
As illustrated in Figure 3, we first process the current and historical frames using a lightweight PoseNet [67, 57] to generate the relative camera poses for photometric reprojection. Subsequently, we utilize these frames to generate both the current feature map $F_{t}$ and historical feature maps $\{F_{t-1},\cdots,F_{t-n}\}$. Following [59, 57], we construct the warped historical features by applying homography warping using the relative camera poses and alternate depth hypothesis planes, which is defined as:
@@ 3.3 Temporal Alignment with Feature Volume Construction
where $\left\{\mathbf{K}_{i}\right\}_{i=0}^{N-1}$ represent the camera intrinsic parameters and $\left\{\left[\mathbf{R}_{0,i}\mid\mathbf{t}_{0,i}\right]\right\}_{i=1}^{N-1}$ denote the extrinsic parameters, respectively. The variable $d_{j}$ represents the hypothesized depth for pixel $\mathbf{p}$ in $F_{t}$. Following this, we aggregate all the warped historical features to create a historical feature volume $\textbf{V}_{tem}^{his}$, which ensures geometric compatibility across varying depth values between the current and historical frames. Next, we lift $F_{t}$ along the depth dimension as described in [57, 72], generating the current feature volume $\textbf{V}_{tem}^{cur}$. By concatenating $\textbf{V}_{tem}^{his}$ with $\textbf{V}_{tem}^{cur}$ following [57], we construct the composite temporal feature volume $\textbf{V}_{tem}$:
@@
The temporal feature volume $\textbf{V}_{tem}$ enhances semantic scene modeling by aligning contextual features across different time steps. In Section 3.4 and Section 3.5, we will detail the methodology for harnessing reliable information through contextual correspondence within $\textbf{V}_{tem}$.
Why Feature Volume Instead of Cost Volume? Conventional temporal depth estimation networks typically build cost volumes by computing the Hadamard product [58, 59] or absolute differences [57] between different feature maps:
[…]
@@
3.4 Cross-frame Pattern Affinity for Relevance Modeling
@@
Similarity Measurement Optimization Analysis. As a widely applied metric in semantic analysis [73, 74, 75] and information retrieval [76, 77], Cosine similarity measures correlations effectively. The cosine similarity between two vectors $\alpha$ and $\beta$ is computed as:
@@
Nevertheless, traditional cosine similarity can yield high similarity scores for vectors that are not truly similar [78]. This issue is acknowledged and addressed through scale-aware isolation [79], which adjusts for variations in pattern scales. However, such methods primarily focus on vector orientations and struggle to assess similarity within densely distributed datasets. To mitigate these limitations, ensemble learning techniques [80], which utilize a diverse array of independent learners, have been employed to improve the accuracy of similarity assessments in dense environments.
@@
3.5 Affinity-based Dynamic Refinement
@@
To this end, we propose to dynamically refine the feature sampling locations based on the identified high-affinity locations and their neighboring relevant regions. The above ideas are implemented using 3D deformable convolutions [81, 82]. Specifically, dynamic refinement is achieved through the introduction of affinity-based correspondence weights and deformable positional offsets. In the context of a sampling grid window $K_{w}$, the formula is expressed as:
@@
3.6 Global Alignment with Unified Transformation
To globally align the geometric context with the temporal context within a unified space for semantic-consistent aggregation, we present the Depth-Hypothesis-Based Transformation (DHBT) as follows. Firstly, the temporal volume $\widetilde{\mathbf{V}}_{{tem}}$ is aligned with the voxel feature volume $\mathbf{V}_{{vox}}$ through depth distribution hypothesis. As depicted in Figure 8, we take the depth hypothesis of $\widetilde{\mathbf{V}}_{{tem}}$ as the distance axis and employ voxel-pooling operation following [30, 18] to splat the volumetric features into the unified space. Following that, we aggregate the voxel feature volume $\mathbf{V}_{{vox}}$ and the temporal feature volume $\widetilde{\mathbf{V}}_{{tem}}$ for reliable information interaction. In the initial stages of training, unregulated temporal information may compromise the learning of voxel features. To address this, we employ a flexible element-wise aggregation strategy:
@@ 3.6 Global Alignment with Unified Transformation
where Zero_Conv is the zero convolution as ControlNet [83] to retain the inherent capabilities of the temporal feature volume. The composed volume $\mathbf{V}_{{com}}$ is then processed through a task-specific head to generate the semantic occupancy voxel or LiDAR semantic labels following previous works [6, 17].
Why Global Alignment Necessary? The Lift-Splat-Shoot (LSS) paradigm is widely employed in bird’s-eye view (BEV) representations, which typically aggregates multi-view image features into a unified space based on depth distributions, enabling the transformation of 2D images into 3D representations [30, 39, 40, 42, 17]. The transformation is performed via the outer product between the 2D image feature of $i^{t}h$ frame and its corresponding depth distribution: $\textbf{F}_{i}^{BEV}=f_{i}^{2d}\otimes d_{i}^{dis}$. The core motivation behind such operation lies in the necessity for constructing a uniform contextual distribution that facilitates more reliable representation and stable learning processes. However, in our framework, the temporal volume $\widetilde{\mathbf{V}}_{{tem}}$ and the voxel feature volume $\mathbf{V}_{{vox}}$ are initially misaligned due to different construction strategies used in their respective representation spaces. Therefore, we employ a unified global transformation approach with the depth distribution hypothesis, to construct reliable volumetric representations.
3.7 Training Objectives
We follow the basic learning objective of MonoScene [16] for semantic occupancy prediction. Standard semantic loss $\mathcal{L}_{{sem}}$ and geometry loss $\mathcal{L}_{{geo}}$ are leveraged for semantic and geometry supervision, while an extra class weighting loss $\mathcal{L}_{ce}$ is also added. To further enforce the ensembled volume, we adopt a binary cross entropy loss $\mathcal{L}_{depth}$ to encourage the sparse depth distribution. The overall learning objective of this framework is formulated as follows:

### v2

@@ III-A1 Preliminary
where $\widehat{\mathbf{V}}$ represents the estimated 3D semantic voxel grid, which aims to approximate the ground truth semantic occupancy or LiDAR semantic labels. For LiDAR semantic segmentation, we use the LiDAR data only for point query to compute evaluation metrics following previous works [huang2023tri, zhang2023occformer].
III-A2 Architectural Design Comparison and Analysis
To estimate high-quality 3D semantic voxel grid $\widehat{\mathbf{V}}$, existing methods attempt to optimize the scene modeling process from geometric or temporal perspectives while neglecting the misalignment issue. Specifically, the geometric modeling solution [zhang2023occformer] leverages naive monocular depth estimation for semantic voxel grid estimation and neglects the uncertainty inherent in the depth estimation process:
@@ III-A2 Architectural Design Comparison and Analysis
where $\otimes$ denotes the outer product. Extr and Mono represent the 2D feature extractor and monocular depth estimator, respectively. Such a process inevitably causes geometric ambiguity when the estimated depth is integrated with the corresponding 2D features. On the other hand, the temporal modeling solution [li2023voxformer] straightforwardly stacks the temporal image frames to construct the semantic voxel grid $\widehat{\mathbf{V}}$, which ignores the positional changes of shared semantic content across different perspectives:
@@
To tackle the misalignment issue, we explore first disentangling the complicated semantic scene comprehension into geometric and temporal context alignment, and further align these contexts globally to compose them together to construct the reliable semantic voxel grid $\widehat{\mathbf{V}}$:
$\widehat{\mathbf{V}}=\texttt{Compose}(\texttt{Geo}(I_{t}^{{rgb}},I_{t-1}^{{rgb}},\ldots),\texttt{Tem}(I_{t}^{{rgb}},I_{t-1}^{{rgb}},\ldots)),$ (4)
where $I_{t}^{{rgb}}$ represents the input temporal RGB image at time step $t$. The functions Geo and Tem denote geometric context extraction and temporal context extraction, respectively, while Compose represents the global fusion of the two contextual representations. In this way, our methods take advantage of the complementary merits of the geometric and temporal representations in a disentangled local-global architecture, which are hierarchically aligned for a reliable context composition.
@@ III-A3 Overall Framework
Geometric Alignment. The voxel feature volume $\mathbf{V}_{{vox}}$ is constructed using a UNet architecture based on a pre-trained EfficientNetB7 [tan2019efficientnet]. The network initially generates features with spatial dimensions of $\mathbb{R}^{H/4\times W/4}$. We then adopt the lifting process following previous studies [philion2020lift, li2023bevdepth, li2024bridging], to form $\mathbf{V}_{{vox}}$ from the contextual information and depth distribution. For depth distribution modeling, we use off-the-shelf monocular [bhat2021adabins] or stereo [cheng2020hierarchical] depth estimation networks. By default, we leverage stereo depth estimation to form our stereo-based pipeline of Hi-SOP(S). Additionally, a monocular-based pipeline of Hi-SOP(M) is presented to enhance versatility for scenarios lacking stereo inputs. To facilitate reliable geometry alignment when constructing $\mathbf{V}_{{vox}}$, we employ a Geometric Confidence-awareness Lifting (GCL) module, which is detailed in Section III-B.
@@
Temporal Alignment. To construct the temporal feature volume $\mathbf{V}_{{tem}}$, we feed current and historical frames into a lightweight PoseNet [guizilini20203d, cai2023riav] to generate temporal feature volume $\mathbf{V}_{{tem}}$ using homography warping. Different from computing matching costs in typical temporal depth estimation methods [long2021multi, watson2021temporal, cai2023riav], we aim to preserve context features within $\mathbf{V}_{{tem}}$. The details are presented in Section III-C.
@@
Following that, we leverage the temporal volume $\mathbf{V}_{{tem}}$ to generate the cross-frame affinity $\hat{\mathbf{A}}$, quantifying contextual relevance/correspondences between current and historical features. This affinity is then used to reassemble the temporal content and dynamically refine the sampling locations, resulting in a reliable aligned temporal volume $\widetilde{\mathbf{V}}_{{tem}}$. Further details on Cross-frame Pattern Affinity (CPA) and Affinity-based Dynamic Refinement (ADR) are presented in Sections III-D and III-E, respectively.
@@
Global Composition. To construct a reliable unified representation with semantically consistency, the temporal feature volume $\widetilde{\mathbf{V}}_{{tem}}$ is aligned with the voxel feature volume $\mathbf{V}_{{vox}}$ in a global view through depth-hypothesis based transformation. These two volumes are composed together to generate the composed volume $\mathbf{V}_{{com}}$. The Depth-Hypothesis-Based Transformation (DHBT) module is detailed in Section III-F.
III-B Geometric Alignment with Confidence-aware Lifting
@@
As introduced in previous works [philion2020lift, li2024bridging], the lifting process establishes volumetric features that store pixel-level context with their associated depth distribution. Nonetheless, [li2024bridging] highlights that this process is inherently ambiguous and prone to producing unreliable representations in challenging regions (e.g., severe occlusion and high reflection) where the depth estimation results are unreliable [li2024bridging]. To facilitate reliable geometry modeling within the voxel feature volume $\mathbf{V}_{{vox}}$, we propose a Geometric Confidence-aware Lifting (GCL) module to explicitly model the geometric information with depth distribution confidence.
@@
As shown in Figure 4, the module takes the depth feature $\textbf{F}_{d}$ from the depth net and the context feature $\textbf{F}_{c}$ from the context net as inputs. To establish pixel-level reliable information for dense prediction, we develop a depth confidence-aware cross-attention mechanism to explicitly indicate the confidence information of the depth distribution and take advantage of the relevant context to complement the low-confidence regions. Specifically, to project $\textbf{F}_{d}$ to a confidence map $\textbf{C}_{d}$ , we first adopt softmaxsoftmax to convert depth cost value $d_{i}$ of $\textbf{F}_{d}$ into a probability form, and then take out the highest probability value among all depth hypothesis planes along the depth dimension as the prediction confidence. The process is formally written as:
@@
where the softmaxsoftmax is applied across the depth dimension and $\mathrm{WTA}$ represents winner-takes-all operation. $D_{max}$ denotes the length of the depth dimension.
Next, we utilize the depth confidence information to enforce the cross-attention for pixel-level reliable geometric modeling. Specifically, we obtain the query $Q_{d}$ from $\textbf{F}_{d}$ by flattening in spatial and depth dimensions following the standard protocol [wang2018non, liao2022wt]. Similarly, the context feature $\textbf{F}_{c}$ is forwarded and its key and value are denoted as $K_{c},V_{c}$, respectively. To reduce computational and memory consumption, we follow [katharopoulos2020transformers, kitaev2020reformer] to compute linear cross-attention:
@@
III-C Temporal Alignment with Feature Volume Construction
The fine-grained nature of the SOP task requires constructing temporally aligned features for accurate and robust perception. Instead of simply stacking input images from various viewpoints as [li2023voxformer], we propose to align the temporal invariant content using an explicit homography transformation.
As illustrated in Figure 3, we first process the current and historical frames using a lightweight PoseNet [guizilini20203d, watson2021temporal] to generate the relative camera poses for photometric reprojection. Subsequently, we utilize these frames to generate both the current feature map $F_{t}$ and historical feature maps $\{F_{t-1},\cdots,F_{t-n}\}$. Following [cai2023riav, watson2021temporal], we construct the warped historical features by applying homography warping using the relative camera poses and alternate depth hypothesis planes, which is defined as:
@@ III-C Temporal Alignment with Feature Volume Construction
where $\left\{\mathbf{K}_{i}\right\}_{i=0}^{N-1}$ represent the camera intrinsic parameters and $\left\{\left[\mathbf{R}_{0,i}\mid\mathbf{t}_{0,i}\right]\right\}_{i=1}^{N-1}$ denote the extrinsic parameters, respectively. The variable $d_{j}$ represents the hypothesized depth for pixel $\mathbf{p}$ in $F_{t}$. Following this, we aggregate all the warped historical features to create a historical feature volume $\textbf{V}_{tem}^{his}$, which ensures geometric compatibility across varying depth values between the current and historical frames. Next, we lift $F_{t}$ along the depth dimension as described in [watson2021temporal, newcombe2011dtam], generating the current feature volume $\textbf{V}_{tem}^{cur}$. By concatenating $\textbf{V}_{tem}^{his}$ with $\textbf{V}_{tem}^{cur}$ following [watson2021temporal], we construct the composite temporal feature volume $\textbf{V}_{tem}$:
@@
The temporal feature volume $\textbf{V}_{tem}$ enhances semantic scene modeling by aligning contextual features across different time steps. In Section III-D and Section III-E, we will detail the methodology for harnessing reliable information through contextual correspondence within $\textbf{V}_{tem}$.
Why Feature Volume Instead of Cost Volume? Conventional temporal depth estimation networks typically build cost volumes by computing the Hadamard product [long2021multi, cai2023riav] or absolute differences [watson2021temporal] between different feature maps:
$\textbf{C(d)}=\frac{1}{N}\sum_{i=1}^{N}\texttt{Match}({f_{0}^{ref},\tilde{f}_{i}^{warp}}),$ (9)
@@
III-D Cross-frame Pattern Affinity for Relevance Modeling
@@
Similarity Measurement Optimization Analysis. As a widely applied metric in semantic analysis [evangelopoulos2013latent, ramachandran2011automated, evangelopoulos2012latent] and information retrieval [rahutomo2012semantic, korenius2007principal], Cosine similarity measures correlations effectively. The cosine similarity between two vectors $\alpha$ and $\beta$ is computed as:
@@
Nevertheless, traditional cosine similarity can yield high similarity scores for vectors that are not truly similar [sarwar2001item]. This issue is acknowledged and addressed through scale-aware isolation [anastasiu2014l2ap], which adjusts for variations in pattern scales. However, such methods primarily focus on vector orientations and struggle to assess similarity within densely distributed datasets. To mitigate these limitations, ensemble learning techniques [xia2015learning], which utilize a diverse array of independent learners, have been employed to improve the accuracy of similarity assessments in dense environments.
@@
III-E Affinity-based Dynamic Refinement
@@
To this end, we propose to dynamically refine the feature sampling locations based on the identified high-affinity locations and their neighboring relevant regions. The above ideas are implemented using 3D deformable convolutions [dai2017deformable, wang2021patchmatchnet]. Specifically, dynamic refinement is achieved through the introduction of affinity-based correspondence weights and deformable positional offsets. In the context of a sampling grid window $K_{w}$, the formula is expressed as:
@@
III-F Global Alignment with Unified Transformation
To globally align the geometric context with the temporal context within a unified space for semantic-consistent aggregation, we present the Depth-Hypothesis-Based Transformation (DHBT) as follows. Firstly, the temporal volume $\widetilde{\mathbf{V}}_{{tem}}$ is aligned with the voxel feature volume $\mathbf{V}_{{vox}}$ through depth distribution hypothesis. As depicted in Figure 8, we take the depth hypothesis of $\widetilde{\mathbf{V}}_{{tem}}$ as the distance axis and employ a voxel-pooling operation following [philion2020lift, li2024bridging] to splat the volumetric features into the unified space. Following that, we aggregate the voxel feature volume $\mathbf{V}_{{vox}}$ and the temporal feature volume $\widetilde{\mathbf{V}}_{{tem}}$ for reliable information interaction. In the initial stages of training, unregulated temporal information may compromise the learning of voxel features. To address this, we employ a flexible element-wise aggregation strategy:
@@ III-F Global Alignment with Unified Transformation
where Zero_Conv is the zero convolution as ControlNet [zhang2023adding] to retain the inherent capabilities of the temporal feature volume. The composed volume $\mathbf{V}_{{com}}$ is then processed through a task-specific head to generate the semantic occupancy voxel or LiDAR semantic labels following previous works [huang2023tri, zhang2023occformer].
Why Global Alignment Necessary? The Lift-Splat-Shoot (LSS) paradigm is widely employed in bird’s-eye view (BEV) representations, which typically aggregates multi-view image features into a unified space based on depth distributions, enabling the transformation of 2D images into 3D representations [huang2021bevdet, li2022bevformer, zhang2023occformer]. The transformation is performed via the outer product between the 2D image feature of $i^{t}h$ frame and its corresponding depth distribution: $\textbf{F}_{i}^{BEV}=f_{i}^{2d}\otimes d_{i}^{dis}$. The core motivation behind such an operation lies in the necessity for constructing a uniform contextual distribution that facilitates more reliable representation and stable learning processes. However, in our framework, the temporal volume $\widetilde{\mathbf{V}}_{{tem}}$ and the voxel feature volume $\mathbf{V}_{{vox}}$ are initially misaligned due to different construction strategies used in their respective representation spaces. Therefore, we employ a unified global transformation approach with the depth distribution hypothesis, to construct reliable volumetric representations.
III-G Training Objectives
We follow the basic learning objective of MonoScene [cao2022monoscene] for semantic occupancy prediction. Standard semantic loss $\mathcal{L}_{{sem}}$ and geometry loss $\mathcal{L}_{{geo}}$ are leveraged for semantic and geometry supervision, while an extra class weighting loss $\mathcal{L}_{ce}$ is also added. To further enforce the ensembled volume, we adopt a binary cross entropy loss $\mathcal{L}_{depth}$ to encourage the sparse depth distribution. The overall learning objective of this framework is formulated as follows: